OpenStudy (anonymous):

in parallelogram PQRS if angle P=5x-8 and angle R=4x+2 then find the degree measure of angle Q

OpenStudy (anonymous):

angle P =angle R 5x-8 =4x+2 5x-4x = 2+8 x =10 angle P=5x-8 =5(10)-8 = 50 -8 angle P=42 angle P +angle Q = 180 angle Q = 180 -42 angle Q = 138If I deserve medal then click good answer